OpenStudy (anonymous):

A fair coin is tossed 8 times. What is the probability of getting at least 6 heads

OpenStudy (kropot72):

This is solved by using the binomial distribution: \[P(6\ heads)=\left(\begin{matrix}8 \\ 6\end{matrix}\right)0.5^{6}0.5^{2}\] \[P(7\ heads)=\left(\begin{matrix}8 \\ 7\end{matrix}\right)0.5^{7}0.5^{1}\] \[P(8\ heads)=\left(\begin{matrix}8 \\ 8\end{matrix}\right)0.5^{8}0.5^{0}\] The required probability is found by adding the above three values of probability.